# Query of CC: Unearthing Large Scale Domain-Specific Knowledge from Public Corpora

### v3

@@ Abstract
Large language models have demonstrated remarkable potential in various tasks, however, there remains a significant scarcity of open-source models and data for specific domains. Previous works have primarily focused on manually specifying resources and collecting high-quality data on specific domains, which significantly consume time and effort. To address this limitation, we propose an efficient data collection method Query of CC based on large language models. This method bootstraps seed information through a large language model and retrieves related data from public corpora. It not only collects knowledge-related data for specific domains but unearths the data with potential reasoning procedures. Through the application of this method, we have curated a high-quality dataset called Knowledge Pile, encompassing four major domains, including stem and humanities sciences, among others. Experimental results demonstrate that Knowledge Pile significantly improves the performance of large language models in mathematical and knowledge-related reasoning ability tests. To facilitate academic sharing, we open-source our dataset and code, providing valuable support to the academic community.
Query of CC: Unearthing Large Scale Domain-Specific Knowledge
from Public Corpora
Zhaoye Fei1,2*, Yunfan Shao1,2♠, Linyang Li1,2♠, Zhiyuan Zeng1,2 Conghui He2, Hang Yan2, Dahua Lin2 and Xipeng Qiu1 ${}^{1}$chool of Computer Science, Fudan University, Shanghai, China ${}^{2}$hanghai AI Laboratory {zyfei20, yfshao19, xpqiu}@fudan.edu.cn {lilinyang, yanhang}@pjlab.org.cn
**footnotetext:  Work done during internship at Shanghai AI Laboratory.footnotetext:  These authors contributed equally to this work.11footnotetext:  Knowledge Pile will be released in https://huggingface.co/datasets/ngc7293/Knowledge_Pile.
@@ 1 Introduction
Large language models (LLMs) are becoming the new trend not only in natural language processing but also in the entire AI community, pioneered by OpenAI ChatGPT and GPT-4 OpenAI (2023). While commercial LLMs are close-sourced, open-source models such as LLaMA Touvron et al. (2023a) and Mistral Jiang et al. (2023) are widely studied by the community since they serve as general base models for building LLM applications. Based on these base models, domain-specific models, show great potential in specific domains, such as medicine Yang et al. (2022); Gao et al. (2023), finance Wu et al. (2023); Zhang and Yang (2023), science Taylor et al. (2022a); Wei et al. (2023), and law Nguyen (2023); Cui et al. (2023). These domain-specific enhanced models are based on specific human-crafted data recipes Azerbayev et al. (2023); Wang et al. (2023c).
@@
However, crafting domain-specific data is very costly. As depicted in Figure 0(a), traditional data collection methods involve the selection of relevant resources by domain experts, followed by data collection and processing by engineers. On the one hand, such endeavors are highly labor-intensive, requiring several months of collaboration between multiple domain experts and engineers for corpus collection. On the other hand, some specific domain-related data distribution may be highly scattered, which poses many challenges for large-scale domain-specific data collection. Therefore, in this paper, we introduce an automatic strategy to retrieve data from public corpora for specific domain knowledge bootstrapping, which we call Query of CC.
In Query of CC, we initially collected seed information in some specific domains, such as keywords, frequently asked questions, and textbooks, to serve as inputs for the Query Bootstrapping stage. Leveraging the great generalization capability of large language models, we can effortlessly expand the initial seed information and extend it to an amount of domain-relevant queries. Inspiration from Wang et al. (2023b) and Xu et al. (2023), we encompassed two stages of expansion, namely Question Extension and Thought Generation, which respectively extend the queries in terms of breadth and depth, for retrieving the domain-related data with a broader scope and deeper thought. Subsequently, based on the queries, we retrieved relevant documents from public corpora, and after performing operations such as duplicate data removal and filtering, we formed the final training dataset.
Otherwise, leveraging Query of CC, we collect a high-quality knowledge dataset Knowledge Pile, which starts from some seed information of four major domains, including STEM, humanities, social sciences, and medical sciences, as well as general knowledge. Utilizing Knowledge Pile, we enhance the Llama and Mistral models through continuing learning. Experimental results indicate that through the Knowledge Pile, both Llama and Mistral enhanced models achieved significant performance improvements over baselines in benchmark tests related to mathematics, knowledge assessments, professional examinations, and some complex reasoning tasks.
@@
  • We propose Query of CC, a self-bootstrapping strategy for automatic data collection to enhance domain-specific knowledge from public corpora.
  • We collect and release Knowledge Pile, based on various categories such as STEM, human science, and social science. The dataset effectively enhances the model’s abilities in mathematical and knowledge-based reasoning.
  • We introduce Llama-QoC and Mistral-QoC models trained on Knowledge Pile, which demonstrate significant improvements on several professional exams and reasoning datasets.
@@ 2.1 Large language model for knowledge-based reasoning
In recent years, significant progress has been made in the field of Natural Language Processing (NLP), driven by the emergence of large language models OpenAI (2023); InternLM-Team (2023); Bai et al. (2023); Sun et al. (2023). Particularly, in the domain of academic and professional examinations, Some language models such as ChatGPT and GPT-4 OpenAI (2023) have demonstrated remarkable success in solving complex tasks, achieving human-like performance through the utilization of the capability of reasoning Wei et al. (2022); Wang et al. (2023a). However, open-source language models lag behind in performance (Like Llama Touvron et al. (2023a), Mistral Jiang et al. (2023) etc.), possibly due to a lack of data.
@@ 2.2 Manual Data Collection
Currently, extensive efforts are being dedicated to the manual collection of specific training data to enhance the capabilities of Large Language Models (LLM) in knowledge-based reasoning. In the field of mathematics, Lewkowycz et al. (2022) undertook the task of gathering approximately 40 billion tokens of data from arXiv and web math pages. They developed a series of Minerva models based on PaLM Chowdhery et al. (2023) and observed that augmenting the model with more mathematical data significantly enhances its proficiency in mathematical reasoning. Similarly, numerous works Azerbayev et al. (2023); Wang et al. (2023c); Paster et al. (2023) have undertaken the collection of mathematics-related data, encompassing papers, web pages, and code, with considerable cost.
In the academic and technological domains, Taylor et al. (2022b) collecting 106 billion academic and technological data. They asserted that the resulting 120B Galactica model surpasses GPT-3 in various academic benchmarks. These works highlight the efficacy of manual data collection in enhancing model performance. However, it is crucial to note that these data collection endeavors are labor-intensive, lack diversity, and present challenges in scalability, thereby posing some constraints on the overall improvement of model performance.
Figure 2: The overview of Query of CC’s two major components: Query Bootstrapping and Data Retrieval.
In contrast to these human-centric collection methods, our approach delves into an automated method for bootstrapping and collecting domain-specific data from public corpora and aims to achieve scalability, efficiency, and cost-effectiveness in data collection.
@@ 2.3 Retrieval-based Data Collection
For large language models, many works Li et al. (2023); Li and Qiu (2023) are based on retrieval to enhance their capabilities. The majority of these focus on retrieving documents relevant to the questions to improve the model’s prior knowledge, thereby enhancing the performance on knowledge-related tasks and reducing hallucinations. Some works Yao et al. (2022) attempt to use retrieval during the training phase for data selection to improve specified downstream tasks. However, the data selected by such methods lacks diversity on the one hand, with the selection scope heavily dependent on the choice of downstream tasks, and on the other hand, it is similar to manually collected data, unable to be expanded in an automated manner. In comparison to these retrieval-based methods, our approach utilizes the bootstrapping method of large models, resulting in more diverse queries for data retrieval, making the training data more diverse and scalable.
3 Query of CC
@@ 3.1 Overview
The overview of Query of CC is illustrated in Figure 2. This framework encompasses two major stages, namely Query Bootstrapping and Data Retrieval. During the Query Bootstrapping phase, the Large Language Model constructs questions and answers centered around the given keywords to retrieve relevant text from public corpora. Through this bootstrapping process, we construct queries around the seed keywords in both depth and breadth, ensuring that the query encompasses a wide range of knowledge and sufficient depth. In the Data Retrieval phase, we employed the BM25 algorithm to retrieve text relevant to specific categories and obtained training data through post-processing steps such as deduplication.
3.2 Query Bootstrapping
To efficiently and comprehensively retrieve high-quality data relevant to the given seed information, we employed the Query Bootstrapping method, inspiration from Wang et al. (2023b). During this phase, we expanded queries around seed information such as category or some related questions, by querying large language models and aims to extend the scope of retrieved data from both breadth and depth perspectives.
@@ 3.2 Query Bootstrapping
To broaden the scope of our queries, we leverage large language models to generate questions relevant to the given seed words which we called Question Extension. Leveraging the capabilities of large models for question generation is evident in expanding the conceptual boundaries of the keywords. This process evolves a narrow category into a more comprehensive representation, encompassing various concepts associated with the given words. Consequently, it significantly enhances the breadth of the query set, ensuring a more comprehensive coverage of various aspects within the target domain. Additionally, to ensure diversity in generation, we set different seeds for LLM inference during experimentation, and employ a higher temperature during the sampling process, thereby ensuring the diversity of the generated questions.
@@
In addition to expanding the range of queries through Question Evolution, we are dedicated to enhancing the depth of thought in generated queries through Thought Generation. In this phase, we are inspired by the work of Wei et al. (2022), employing large-scale language models to produce the cognitive processes necessary for answering questions. This enables us to acquire detailed and insightful responses. This approach supports a more thorough exploration of concepts related to seed information and captures the cognitive processes essential for answering questions. The generated thoughts not only serve for retrieving seed information-related data but also, by expanding queries, offer additional context and relevant knowledge, contributing to a more comprehensive understanding of the subject matter.
Query post processing
We conducted post-processing operations on queries, involving two main stages: pre-cleaning before entering the seed information pool and cleaning and filtering before entering the query pool. On one hand, the generated questions underwent a re-cleaning process and were incorporated into the seed information pool for the next rounds of question bootstrapping and thought generation. This cleaning step involved removing incompletely generated language data to prevent the influence of non-natural language. On the other hand, before entering the query pool, we employed Minhash-LSH deduplication on queries to mitigate performance inefficiencies resulting from redundant retrievals.
@@ 3.3 Data Retrieval
Based on the query bootstrapping stage, we get extensive and in-depth queries. During the data retrieval stage, utilizing the enriched queries, we employ the BM25 Robertson and Walker (1994) algorithm to retrieve data from general public corpora. BM25 is a widely adopted relevance calculation method commonly used by search engines. It calculates the relevance score between the given query and target documents by weighting and summing the matching degree of keywords in the query with the target documents. Efficiency is the reason why we use BM25 to calculate the relevance. When dealing with billion data, performing relevance calculations for each query against every document becomes exceedingly challenging, while BM25 rapidly retrieves documents relevant to the target query. Compared with Dense Retriever Karpukhin et al. (2020), it may incur a potential loss in retrieval accuracy, but the latter comes with an unbearable high computational cost. Exploring the potential impact of retriever selection on the quality of collected data might be a pivotal direction for future research.
@@
3.4 Data Post-processing
To enhance the quality of the retrieved dataset, we conducted a series of filtering procedures. Taking inspiration from the filtering methods proposed by  Wenzek et al. (2020), we trained a scoring model to detect high-quality data. This model was trained by the selected high-quality data as positive samples and using the randomly sampled instances from Common Crawl as negative examples. we used this model to score all the retrieved data, and a specified threshold was applied to filter the data. Also, we filter the non-English documents and the code in retrieved data.
Additionally, considering the detrimental impact of data duplication on model training, we employed the Minhash-LSH and n-gram methods to deduplicate the retrieved data. Duplicate data with a similarity threshold of 0.8 was removed from the dataset.
@@ 3.4 Data Post-processing
Proof of Pile
Mathematics reasoning
human collection
arXiv, Textbooks, Lib., Stack Exchange, ProofWiki, MATH
8.3B
OpenWebMath
Mathematics reasoning
human collection Common Crawl 14.7B
MATHPILE
Mathematics reasoning
human collection
arXiv, Textbooks, Lib., Stack Exchange, ProofWiki, MATH,Web
9.5B
AcademicGPT
Academic
human collection
arXiv, Unpaywall,Top Universities, Pubmed, Common Crawl, Semantic Scholar, Wiki
370B
Galactica
Academic
human collection
Papers, Code, Reference Material, Knowledge Bases, Common Crawl, Prompts
106B
Knowledge Pile (ours)
Mathematics reasoning & Knowledge
automatic query and retrieve
Public Corpora
188B
Table 1: Comparation of Knowledge Pilewith other specific domain knowledge dataset. In this table, most data scales are derived from publicly released research papers, while the data scale for the Knowledge Pile is obtained through tokenized data analysis using the Llama2 tokenizer.
4 Knowledge Pile
Leveraging Query of CC, Based on queries of some knowledge categories, we retrieved high-quality knowledge data from processed public corpora. We call the collected datasets as Knowledge Pile. In this section, we will introduce the analysis of queries (Section 4.1) and the statistic information of Knowledge Pile (Section 4.2), See more implement details in Appendix C.

### v4

# Unearthing Large Scale Domain-Specific Knowledge from Public Corpora

@@ Abstract
Large language models (LLMs) have demonstrated remarkable potential in various tasks, however, there remains a significant lack of open-source models and data for specific domains. Previous work has primarily focused on manually specifying resources and collecting high-quality data for specific domains, which is extremely time-consuming and labor-intensive. To address this limitation, we introduce large models into the data collection pipeline to guide the generation of domain-specific information and retrieve relevant data from Common Crawl (CC), a large public corpus. We refer to this approach as Retrieve-from-CC. It not only collects data related to domain-specific knowledge but also mines the data containing potential reasoning procedures from the public corpus. By applying this method, we have collected a knowledge domain-related dataset named $\mathsf{Retrieve}$-$\mathsf{Pile}$, which covers four main domains, including the sciences, humanities, and other categories. Through the analysis of $\mathsf{Retrieve}$-$\mathsf{Pile}$, Retrieve-from-CC can effectively retrieve relevant data from the covered knowledge domains and significantly improve the performance in tests of mathematical and knowledge-related reasoning abilities. We have released $\mathsf{Retrieve}$-$\mathsf{Pile}$ at https://huggingface.co/datasets/Query-of-CC/Retrieve-Pile.
@@ 1 Introduction
Large language models (LLMs) are becoming the new trend not only in natural language processing but also in the entire AI community, pioneered by OpenAI ChatGPT and GPT-4 (OpenAI, 2023). While commercial LLMs are close-sourced, open-source models such as LLaMA (Touvron et al., 2023a) and Mistral (Jiang et al., 2023) are widely studied by the community since they serve as general base models for building LLM applications. Based on these base models, domain-specific models, show great potential in specific domains, such as medicine (Yang et al., 2022; Gao et al., 2023), finance (Wu et al., 2023; Zhang & Yang, 2023), science (Taylor et al., 2022; Wei et al., 2023), and law (Nguyen, 2023; Cui et al., 2023). These domain-specific enhanced models are based on specific human-collected dataset (Azerbayev et al., 2024; Wang et al., 2023c).
However, crafting domain-specific data is very costly. As depicted in Figure 1(a), traditional data collection methods involve the selection of relevant resources by domain experts, followed by data collection and processing by engineers. On the one hand, such endeavors are highly labor-intensive, requiring several months of collaboration between multiple domain experts and engineers for corpus collection. On the other hand, some specific domain-related data distribution may be highly scattered, which poses many challenges for large-scale domain-specific data collection. Therefore, in this paper, we introduce an automatic strategy to retrieve data from public corpora for specific domain knowledge, which we call Retrieve-from-CC.
In Retrieve-from-CC, we initially collected seed information in some specific domains, such as keywords, frequently asked questions, and textbooks, to serve as inputs for the Query Expanding stage. Leveraging the great generalization capability of LLMs, we can effortlessly expand the initial seed information and extend it to an amount of domain-relevant queries. Inspiration from Wang et al. (2023b) and (Xu et al., 2024), we encompassed two stages of expansion, namely Question Extension and Thought Generation, which respectively extend the queries in terms of breadth and depth, for retrieving the domain-related data with a broader scope and deeper thought. Subsequently, based on the queries, we retrieved relevant documents from public corpora, and after performing operations such as duplicate data removal and filtering, we formed the final training dataset.
@@
Leveraging Retrieve-from-CC, we collect a high-quality knowledge dataset $\mathsf{Retrieve}$-$\mathsf{Pile}$, which starts from some seed information of four major domains, including STEM, humanities, social sciences, and medical sciences, as well as general knowledge. Utilizing $\mathsf{Retrieve}$-$\mathsf{Pile}$, we enhance the Llama and Mistral models through further pre-training. Experimental results indicate that through the $\mathsf{Retrieve}$-$\mathsf{Pile}$, Llama and Mistral enhanced models achieved significant performance improvements over baselines in benchmark tests related to mathematics, knowledge assessments, professional examinations, and some complex reasoning tasks.
@@
  • We propose Retrieve-from-CC, a data collection pipeline to retrieve domain-specific knowledge from public corpora, which introduces LLMs to extend query and retrieve domain-related data from public corpora.
  • We collect and release a knowledge-related corpora $\mathsf{Retrieve}$-$\mathsf{Pile}$ based on Retrieve-from-CC from Common Crawl, a large-scale public corpora, which includes various categories such as STEM, human science, and social science.
  • We have analysed the quality and statistical properties of $\mathsf{Retrieve}$-$\mathsf{Pile}$. We examined the distribution of web domain to show the performance of Retrieve-from-CC in collecting scattered information, then, we compared the educational value of $\mathsf{Retrieve}$-$\mathsf{Pile}$  with that of other Open-source knowledge-related datasets, to demonstrate the high educational value of our dataset.
  • We train several language models on $\mathsf{Retrieve}$-$\mathsf{Pile}$, which demonstrate significant improvements on several professional exams and reasoning datasets.
@@ 2.1 Large language model for knowledge-based reasoning
In recent years, significant progress has been made in the field of Natural Language Processing (NLP), driven by the emergence of large language models (OpenAI, 2023; InternLM-Team, 2023; Bai et al., 2023; Sun et al., 2023). Particularly, in the domain of academic and professional examinations, Some language models such as ChatGPT and GPT-4 (OpenAI, 2023) have demonstrated remarkable success in solving complex tasks, achieving human-like performance through the utilization of the capability of reasoning (Wei et al., 2022; Wang et al., 2023a). However, open-source LLMs lag in performance (Like Llama (Touvron et al., 2023a), Mistral (Jiang et al., 2023) etc.), possibly due to a lack of data.
@@ 2.2 Manual Data Collection
Extensive efforts are being dedicated to the manually collection of specific training data to enhance the capabilities of Large Language Models (LLM) in knowledge-based reasoning. In the field of mathematics, Lewkowycz et al. (2022) undertook the task of gathering approximately 40 billion tokens of data from arXiv and web math pages. They developed a series of Minerva models based on PaLM (Chowdhery et al., 2023) and observed that augmenting the model with more mathematical data significantly enhanced its proficiency in mathematical reasoning. Similarly, numerous works (Azerbayev et al., 2024; Wang et al., 2023c; Paster et al., 2024) have undertaken the collection of mathematics-related data, including papers, web pages, and code, at considerable cost.
In the academic and technological domains, Taylor et al. (2022) collected 106 billion tokens of academic and technological data. They asserted that the resulting 120B Galactica model surpasses GPT-3 on various academic benchmarks. These works highlight the effectiveness of manual data collection in enhancing model performance. However, it is crucial to note that these data collection endeavors are labor-intensive, and present scalability challenges, thereby posing some constraints on the overall improvement of model performance.
Figure 2: The overview of Retrieve-from-CC’s two major components: Query Expanding and Data Retrieval.
Overall, manually curated domain-specific datasets typically require substantial human labor for data collection and filtering. For example, the Pile dataset contains 22 distinct web domains, each requiring significant human input for collection, formatting, and initialization. The OpenWebText dataset utilizes multiple filters designed to extract high-quality, domain-specific text. In contrast, the Retrieve-from-CC collects high-quality domain-specific data by simply providing relevant keywords, thereby minimizing the need for manual intervention, reducing human effort required to collect domain-specific data.
@@ 2.3 Retrieval-based Data Collection
Many works (Li et al., 2023; Li & Qiu, 2023) utilized retrieval methods to enhance their capabilities. The majority of these focus on retrieving documents relevant to the questions to improve the model’s prior knowledge, thereby enhancing the performance on knowledge-related tasks and reducing hallucinations. Also, (Yue et al., 2024) retrieved related data for enhancing the instruction synthetic. For data collection, some works (Yao et al., 2022) attempt to use retrieval during the training phase for data collection to improve specified downstream tasks. However, Retrieving specified information for specific downstream tasks relies on the data of those tasks, making it difficult to automate and scale. In contrast, our method introduces LLMs to automatically extend domain-related queries, which enhances the automation and scalability of data collection.
3 Retrieve-from-CC
@@ 3.1 Overview
To collect domain-specific data at a lower cost, we propose a retrieval-based method for data collection, which utilizes LLMs to expand keywords into a variety of domain-related queries. These queries are then used to retrieve relevant data from public corpora, the process we refer to as Retrieve-from-CC. An overview of Retrieve-from-CC is illustrated in Figure 2. This framework consists of two main stages: Query Expanding and Data Retrieval. The input to the entire pipeline consists of seed keywords related to the target domain. These seed keywords are expanded into a broader set of domain-relevant queries using LLMs during the Query Expanding stage. In the Data Retrieval stage, the generated queries is used as queries to retrieve relevant data from public corpora.
As shown in Figure 2, during the Query Expanding stage, LLMs generate questions and answers centered around the provided keywords. Both the generated questions and the corresponding answers serve as domain-relevant queries that will be used in the Data Retrieval phase. During the Data Retrieval phase, we employ the BM25 algorithm to retrieve documents relevant to the queries, and then we obtain the final data.
3.2 Query Expanding
To efficiently and comprehensively retrieve high-quality data relevant to the given seed information, we employed the Query Expanding, inspired by Wang et al. (2023b). This phase consists of two steps aimed at broadening the scope of the final query. First, we utilize LLMs to generate questions related to the given keywords or other seed information, we refer it as ’Question Extension.’ Subsequently, using these questions, we prompt the LLMs to provide answers and generate reasoning strategies for solving the questions, thereby retrieving thought-related data from public corpora.
@@ 3.2 Query Expanding
To expand the scope of our queries, we utilize LLMs to generate questions relevant to the provided seed words. By leveraging the generalization capabilities of LLMs, we can easily generate a series of questions related to the seed information, thereby expanding the conceptual boundaries of the target domain. Figure 2 illustrates an example of ’Question Extension’. Given the seed word ’biology,’ LLMs generate related questions, such as those concerning ’genetics’ and ’meiosis’. This process evolves limited seed information into a more comprehensive representation that encompasses a range of related concepts. Consequently, this approach significantly enhances the breadth of the query set, ensuring more comprehensive coverage of various aspects within the target domain. All generated questions are saved in the seed information pool for future iterations and also serve as queries for data retrieval.
@@
In addition to expanding the range of queries through Question Extension, we also notice that the answer and the reasoning path are important for ensuring quality. For the generated questions in ’Question Extension’, we employ LLMs to generate answers and the reasoning processes required to obtain them. This enables us to acquire detailed and insightful responses. This approach supports a more thorough exploration of concepts related to seed information and facilitates the generation of cognitive processes essential for answering questions. These reasoning processes are also used as queries for data retrieval. While we found that some of these thought data may contain errors or grammatical issues, we still use them as queries to retrieve accurate data from public corpora.
Post processing
After generation (both in ’Question Extension’ and ’Thought Generation’), the generated data is stored in a seed information pool for the next iteration and also in a query pool for data retrieval. Both processes involve the same post-processing methods: cleaning and deduplication. The cleaning step removes incomplete language data to eliminate the impact of non-natural language. For the deduplication stage, Minhash-LSH (Broder, 1997) was employed to remove duplicate data.
@@ 3.3 Data Retrieval
Based on the query expanding stage, we get extensive and in-depth queries. During the data retrieval stage, utilizing the enriched queries, we employ the BM25 (Robertson & Walker, 1994) algorithm to retrieve data from general public corpora. BM25 is a widely adopted relevance calculation method commonly used by search engines. It calculates the relevance score between the given query and target documents by weighting and summing the matching degree of keywords in the query with the target documents. Efficiency is the reason why we use BM25 to calculate the relevance. When dealing with billion data, performing relevance calculations for each query against every document becomes exceedingly challenging, while BM25 rapidly retrieves documents relevant to the target query. Compared with Dense Retriever (Karpukhin et al., 2020), it may incur a potential loss in retrieval accuracy, but the latter comes with an unbearable high computational cost. Exploring the potential impact of retriever selection on the quality of collected data might be a valuable direction for future research.
@@
4 $\mathsf{Retrieve}$-$\mathsf{Pile}$
@@
Proof of Pile Mathematics reasoning human collection arXiv, Textbooks, Lib., Stack Exchange, ProofWiki, MATH 8.3B
OpenWebMath Mathematics reasoning human collection Common Crawl 14.7B
MATHPILE Mathematics reasoning human collection arXiv, Textbooks, Lib., Stack Exchange, ProofWiki, MATH,Web 9.5B
AcademicGPT Academic human collection arXiv, Unpaywall,Top Universities, Pubmed, Common Crawl, Semantic Scholar, Wiki 370B
Galactica Academic human collection Papers, Code, Reference Material, Knowledge Bases, Common Crawl, Prompts 106B
$\mathsf{Retrieve}$-$\mathsf{Pile}$ (ours) Mathematics reasoning & Knowledge automatic query and retrieve Public Corpora 188B
Table 1: Comparation of $\mathsf{Retrieve}$-$\mathsf{Pile}$with other specific domain knowledge dataset. In this table, most data scales are derived from publicly released research papers, while the data scale for the $\mathsf{Retrieve}$-$\mathsf{Pile}$ is obtained through tokenized data analysis using the Llama2 tokenizer.
Leveraging Retrieve-from-CC, Based on queries of some knowledge categories, we retrieved several knowledge-related data from processed public corpora. We call the collected datasets as $\mathsf{Retrieve}$-$\mathsf{Pile}$. In this section, we will introduce the analysis of queries (Section 4.1) and the analysis of $\mathsf{Retrieve}$-$\mathsf{Pile}$ (Section 4.2). Also, we train several language models to show the improvement of $\mathsf{Retrieve}$-$\mathsf{Pile}$ in some knowledge-related reasoning benchmarks (Section 4.3). Otherwise, we discuss about the different when improving different language models using $\mathsf{Retrieve}$-$\mathsf{Pile}$ (Section 4.3). See more implementation details in the Appendix.